OpenStudy (hannahduhe):

find the lengths of the missing sides in the triangle. Write your answers as integers or as decimals rounded to the nearest tenth LOOK IN COMMENTS FOR TRIANGLE

OpenStudy (hannahduhe):



OpenStudy (anthonyym):

Do you know sine, cosine, tangent? http://www.mathwords.com/s/sohcahtoa.htm

OpenStudy (anthonyym):

Or do you know the special 45-45-90 triangle?

OpenStudy (hannahduhe):

yes

OpenStudy (hannahduhe):

not the special 45-45-90

OpenStudy (anthonyym):

So to find X what would you use? sine cosine or tangent

OpenStudy (hannahduhe):

sine

OpenStudy (anthonyym):

Correct Remember sin(angle) = opposite / hypotenuse Now try plugging in your given info to get an equation to solve for the hypotenuse

OpenStudy (hannahduhe):

would the equation be \[\sin(45)\frac{ 7}{ x }\]

OpenStudy (anthonyym):

Close, you need the equals sign. sin(45) = 7/x

OpenStudy (anthonyym):

Now rearrange to solve for x.

OpenStudy (hannahduhe):

\[7\times \sin (45)\]

OpenStudy (anthonyym):

No, try cross multiplying. You know what that is?

OpenStudy (hannahduhe):

kinda

OpenStudy (hannahduhe):

so how would i set the equation up

OpenStudy (anthonyym):

\[\sin(45) = \frac{ 7 }{ x }\] x*sin(45) = 7